OpenStudy (anonymous):

From 1970 to 1990, the average cost C of a new car (in dollars) can be approximated by the model C = 34.3t^2+4086 where t is the year, with t = 0 corresponding to 1970. During which year was the average cost of a new car $10000?

OpenStudy (campbell_st):

let C = 10000 so 10000 = 34.3t^2 + 4086 subtract 4086 from booth sides 5914 = 34.3t^2 t^2 = 5914/34.3 \[t = \pm \sqrt{5914/34.3}\]

OpenStudy (campbell_st):

you'll need to evaluate for t then add it to 1970 to find the year

OpenStudy (anonymous):

so the answer would be.....

OpenStudy (campbell_st):

only consider the positive t value

OpenStudy (campbell_st):

its written above ... just pick up the calculator to evaluate t... then add it to 1970